OpenStudy (anonymous):

Graph the following piece wise function. State the values for f(0) and f(-2). Please show all of your work and submit graph. f(x)={x^2-3 if x>-2@ 3 if-5 ≤x< -2@-x+3 if x< -5)

OpenStudy (anonymous):

Piece wise functions are very simple. You just evaluate f(x) for a certain function depending on what range x falls into. \[\Large f(x) = \left\{ {\begin{array}{*{20}{c}} {x > - 2}&{}&{{x^2} - 3}\\ { - 5 \le x \le - 2}&{}&3\\ {x < - 5}&{}&{ - x + 3} \end{array}} \right.\] Just start an arbitrary x value to one side of the piecewise function, and increment (or in this case decrement) x by 1 re-evaluating at each point. So when... x = -1, f(x) = x^2 -3 = (-1)^2 - 3 = -2, so plot a point at (-1,-2). x = -2, f(x) = 3, plot (-2,3) x = -3, f(x) = 3, plot (-3,3) x = -4, f(x) = 3, plot (-4,3) x = -5, f(x) = 3, plot (-5,3) x = -6, f(x) = -x+3 = -(-6) + 3 = 0, plot (-6,9) And so on.

OpenStudy (anonymous):

is this the answer that I should use?

OpenStudy (anonymous):

can you show the graph?!

OpenStudy (anonymous):

No, because it's REAL hassle to try and get a graph on here. Can't you just plot the graph yourself?

OpenStudy (anonymous):

yes but what is the input so I can plot it?

OpenStudy (anonymous):

I havent quite figured out how to input it into wolphram alpha

OpenStudy (anonymous):

Well... the function is piece wise, so it depends on what range the input falls into. There's three functions to plot: f(x) = x^2 - 3, f(x) = 3 and f(x) = -x + 3 but at which point you plot them depends on what the value of x is at that point. I really can't explain it any better than that.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

im just wondering how to type that into the calculator